OpenStudy (anonymous):

if the problem was 1^2+9^2=c^2 then the answer would be radical 82, right?

myininaya (myininaya):

1+81=c^2 82=c^2 are you finding the side of a right triangle? or you finding some kind of length, because it would also be -sqrt(82)

OpenStudy (anonymous):

yes. although radical does not always mean square root.

myininaya (myininaya):

yes dhatraditya is right

myininaya (myininaya):

but you have two answers unless you are looking for a measurement

OpenStudy (anonymous):

well, i'm finding the length of the unknown side of the right triangle, which i guess is c

myininaya (myininaya):

ok so it is sqrt(82) gj

OpenStudy (anonymous):

then it will be the positive square root, since length cannot be negative.

OpenStudy (anonymous):

so i was right, right?

myininaya (myininaya):

it is sqrt(82)

OpenStudy (anonymous):

thank goodness!